OpenStudy (anonymous):

Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 11 3x + 4y = 24 Any help is appreciated...

OpenStudy (anonymous):

multiply 2 with eq:2 and 3 with eq: 1 6x+3y=33 6x+8y=48 now subtract eq: 1 from eq: 2 -5y=-15 or 5y=15 y=15/5 y=3 now put y=3 in eq:1 or eq:2 6x+3(3)=33 6x+9=33 6x=33-9 6x=24 x=24/6 x=4

OpenStudy (anonymous):

its a simultaneous equation like if there are 2 boys saving money one of them gets 6$each week and another gets 5$ and first and next week he gets 0.25 extra and so on so at which point thier money would be same

OpenStudy (anonymous):

My solution was right or wrong dear?

OpenStudy (anonymous):

right it is is elimination method i just wrote the scenario